Stock R has a beta of 2.0, Stock S has a beta of 0
Answer #1

Hello,

we can find the required return by using the formula

Ke = Rf + B (Rm - Rf), where Rf is risk free rate of return, B is beta, Rm is the market rate of return or average return.

Required return on riskier stock = 5 + 2 (10-5) = 15%

Required return on less risky stock = 5 + 0.45 (10-5) = 7.5%

Therefore the difference between riskier stock and less risky stock is 15% - 7.5% = 7.5%.

Thankyou.